Bill walk north for 251 meters and then turns go 16Degrees North of West. Find the magnitude and distance of the resultant vector 


1
Expert's answer
2020-10-21T09:52:25-0400
"\vec{s}=(0,251)+(251\cos{16},251\sin{16})=(241.3,320.2)\ m"

"s=\sqrt{241.3^2+320.2^2}=\ 401\ m"

"\theta=\arctan{\frac{241.3}{320.2}}=37\degree\text{W of N}"